How many 3/8 - lb bags of trail mix can be made from 6 3/8 lbs of ingredients?
Found 2 solutions by Alan3354, addingup:
Answer by Alan3354(69443)

How many 3/8 - lb bags of trail mix can be made from 6 3/8 lbs of ingredients?
=============
Divide 6 3/8 by 3/8.

Answer by addingup(3677)

(6 3/8)/3/8 = (51/8)/3/8 = 51/8*8/3 = 408/24 = 17 bags